ortad? (b) If the cally onto a floor, hitting with a speed of 24 m/s. It rebounds with an initial speed of 9.8 m/s. (a) What impulse acts on the ball during the contact? (b) i ball is in contact with the floor for 0.0357 s, what is the magnitude of the average force on the floor from the ball? (a) Number 22 72 Units Tkg-m/s or Ns (b) Number 636.4
Answer #1

(a) Impulse,

              I = F.Δt = Δp = m(v - u)

                  =1.6(24 - (- 9.8)) = 19.52 kg.m/s

(b) Δt = 0.0357 s

Average force,

             F = I/Δt = 19.52/0.0357 = 546.77 N